-8(2×2)+3×7+(-2)


1
Expert's answer
2022-06-13T15:55:51-0400

At first we have to do multiplication in parentheses::

"2\u00d72=4,"

so we have:

"-8(2\u00d72)+3\u00d77+(-2)=-8\u00d74+3\u00d77+(-2)."

The next we have to do 2 multiplications:

"-8\u00d74=-32,"

"3\u00d77=21."

So:

"-8\u00d74+3\u00d77+(-2)=-32+21+(-2)="

"=-(32-21)+(-2)=-11+(-2)=-13."


Answer: -13.